How much water must be added to 180 grams of a solution that is 35% acetic acid to produce a solution that is 10% acetic acid?
Answer by stanbon(57387)

Let the amount of water added be "x".
Start: 0.65(180 grs.)= 117 grs. of water
Add: x grams of water
End up with: 0.9(180+x) grs of water
EQUATION:
water + water = water
117+x= 0.9(180+x)
117+x=162+0.9x
0.10x=45
x= 450 grams of water (amount of water added)
Cheers,
Stan H.